OpenStudy (anonymous):

During January 50,000 new vehicles were sold in Ontario. About 20% of these were leased. How many vehicles were leased?

OpenStudy (anonymous):

you have to find 20% of 50000

OpenStudy (anonymous):

= 20/100 * 50000

OpenStudy (anonymous):

Thanks Cause 20/100=0.2 and then 0.2* 50,000=10,000 It now all makes sense